# CuClarabel: GPU Acceleration for a Conic Optimization Solver

### v2

@@ 3.2 Mixed parallel computing strategy
The cone operations related to steps 3, 5, 7, and 8 above require only information local to each constituent cone, and hence can be executed concurrently with respect to individual cones. We propose the mixed parallel computing strategy for a cone operation across different types of cones. Each family of cones is handled in parallel, and families of cones are handled sequentially.
As stated earlier, the cone operations related to zero cones and nonnegative cones are simply vector additions and multiplications, which have already been parallelized in CUDA. In our implementation we aggregate all zero cones into a single zero cone in our preprocessing step. The same holds for nonnegative cones. For the remaining cones, we parallelize within each family of cones (i.e., second-order cone, exponential cone, etc.). For each family of cones, we allocate a thread to each cone belonging to that family, and execute in parallel. We synchronize threads after executing for each type of cone. This adheres to the SIMD computing paradigm, as each type of cone has its own set of instructions for updating its scaling matrix.
Note that the SIMD GPU computing structure naturally favors balanced workloads, i.e., the workloads in each thread should be similar so that the synchronization will not take too much time. The exponential and power cones are 3-dimensional nonsymmetric cones that all cone operations have balanced workload among the same class of cones, thus we can parallelize the computation for each cone. On the other hand, the second-order cones may vary in dimensionality, which may not mesh well with the SIMD computing paradigm. In most use cases for large-scale second-order cone problems, e.g. optimal power flow problems [2] and finite-element problems [5], the second-order cones are of small (less than 5) dimensionality; thus the effect of this workload imbalance is negligible. Regardless, we support second-order cones of all sizes. In the next section, we outline a preprocessing procedure for our second-order cones we use to balance workloads.
We illustrate our mixed parallel computing strategy via step 3, the scaling matrix update. Recall that the conic constraint $\mathcal{K}$ can be decomposed as a Cartesian product of $p$ constituent atomic cones $\mathcal{K}_{1}\times\dots\times\mathcal{K}_{p}$ that are ordered by cone types. We assume $\mathcal{K}_{1}$ is a zero cone, $\mathcal{K}_{2}$ is a nonnegative cone, $\mathcal{K}_{3}$ to $\mathcal{K}_{i}$ are second-order cones, $\mathcal{K}_{i+1}$ to $\mathcal{K}_{j}$ are exponential cones and $\mathcal{K}_{j+1}$ to $\mathcal{K}_{p}$ are power cones. Due to our composition into constituent atomic cones, $H$ is a block-diagonal matrix,
@@
For the update of $H$ at iteration $k+1$, we first set diagonal terms of the scaling matrix $H_{1}^{k+1}$ for the zero cone to all 00s, and then update the diagonals of the scaling matrix $H_{2}^{k+1}$ for the nonnegative cone by element-wise vector division, as in the Nesterov-Todd (NT) scaling [51]. These computations are already parallelized by CUDA. For conic constraints other than the zero cone and the nonnegative cone, the corresponding parts in the scaling matrix $H$ are no longer diagonal, but we observe $H_{t}$ only requires local information within each constituent cone, and thus we can solve for each $H_{t}$ concurrently, independent to the other cones. We implement kernel functions _soc_update_H(), _exp_update_H() and _pow_update_H() for second-order cones, exponential cones and power cones respectively. The kernel function will process one cone operation per thread, and the kernel functions belongs to the same class of cones are executed in parallel and finally synchronized by calling @sync. Details regarding the update functions for each type of cone and Algorithm 1 illustrating the update of scaling matrix $H$ can also be found in §C.
@@
3.3 Second-order cone preprocessing for balanced workloads
We can decompose any large second-order cone into a set of small second-order cones of fixed dimensionality $d$. The workload is then balanced for each second-order cone, adhering to the SIMD computing paradigm. For an $n$-dimensional second-order cone $\mathcal{K}_{soc}^{n}\{x\in\textbf{R}^{n}\mid x_{1}\geq\sqrt{\sum_{i=2}^{n}x_{i}^{2}}\}$, we can decompose it into the equivalent set of $q=\lceil(n-2)/(d-2)\rceil$ second-order cones:
$\begin{array}[]{cl}x_{1}&\geq\sqrt{x_{2}^{2}+\cdots+x_{d-1}^{2}+z_{1}^{2}}\\ z_{1}&\geq\sqrt{x_{d}^{2}+\cdots+x_{2d-2}^{2}+z_{2}^{2}}\\ &\vdots\\ z_{q}&\geq\sqrt{x_{n-d}^{2}+\cdots+x_{n}^{2}}\end{array},$
where $z\in\textbf{R}^{q}$ is an a slack variable. Each of the first $q-1$ second-order cones are $d$-dimensional, while the $g^{\textrm{th}}$ cone has dimension not larger than $d$. The workload is now balanced among these $p$ cones.
@@ 3.5 Solving linear systems
Most of the computation time for our interior point method is spent in factorizing the matrix $K$ and in the three backsolve operations in (7). Our GPU implementation also leverages the power of the newly released sparse linear system solver CUDSS [36] for the $LDL^{T}$ factorization and backsolve operations. The iterative refinement [52] is implemented to increase the numerical stability of the backsolve operation.
@@
    Solve for the residual at step $i$: $r_{i}=b-Kx_{i-1}$, where $x_{i-1}$ is our guess for iteration $i-1$. (Full precision)
@@
In all our matrix solves, given matrix $K$ and right-hand side $b$, we factorize the lower precision copy of $K$, and then solve the linear system in this lower precision. A regularization parameter $\delta$ is added to $K$ to increase the numerical stability of matrix factorization. It is the sum of a static regularization and a dynamic regularization. To offset the regularization effect and rounding errors from the lower precision in backsolves, we solve in full precision for the other steps, i.e., we compute the residual $r_{i}$ and save the update $\{x_{i}\}$ in full precision. We loop until either (1) we reach a pre-specified number of max steps for iterative refinement, or (2) the $\ell_{\infty}$ norm of $b-Kx_{i}$ reaches a certain threshold, specifically $\left\|b-Kx_{i}\right\|_{\infty}\leq t_{\text{abs}}+t_{\text{rel}}\left\|b\right\|_{\infty}$. For the mixed precision, we set the static regularization to the square root of machine precision, i.e., $\sqrt{\epsilon}$, and take a maximum of $10$ iterative refinement steps. For the full-precision solve, $t_{\text{abs}}$ (absolute tolerance) and $t_{\text{rel}}$ (relative tolerance) are both $10^{-12}$ by default.
Although the mixed precision for the iterative refinement can improve the numerical stability for solving linear system in lower precision, it is to be expected that the mixed precision may take longer time to converge or fail in cases where the matrix $K$ is extremely ill-conditioned. Caution is required when using mixed precision for numerically hard problems, e.g., conic programs with exponential cones.
@@ 4 Numerical experiments
We have benchmarked our Julia GPU implementation of the Clarabel solver [^1] https://github.com/cvxgrp/CuClarabel against the state-of-the-art commercial interior-point solvers Mosek [12] and Gurobi [53]. We also include our Rust CPU implementation of Clarabel solver with the 3rd-party multithreaded supernodal LDL factorization method in the faer-rs package [54]. We have included benchmark results for several classes of problems including quadratic programming (QP), second-order cone programming (SOCP) and exponential cone programming. All benchmarks are performed using the default settings for each solver, with pre-solve disabled where applicable to ensure equivalent problem-solving conditions. No additional iteration limits are imposed beyond each solver’s internal defaults. All experiments were carried out on a workstation with Intel(R) Xeon(R) w9-3475X CPU @ 4.8 GHz with 256 GB RAM and NVIDIA GeForce RTX 4090 24GB GPU. All benchmarks tests are scripted in Julia and access solver interfaces via JuMP [82]. We use Rust compiler version 1.76.0 and Julia version 1.10.2.
@@ 4.1 Benchmarking metrics
We choose the same benchmarking tests as used in the Clarabel solver [15], and compare our results using metrics that are commonly used when comparing solve time across different solvers. For a set of $N$ test problems, we define the shifted geometric mean $g_{s}$ as
@@
where $\mathcal{I}_{\leq\tau}(u)=1$ if $\tau\leq u$ and $\mathcal{I}_{\leq\tau}(u)=0$ otherwise. We also compute the absolute performance profile $f^{a}_{s}:\textbf{R}_{+}\mapsto[0,1]$, which denotes the fraction of problems solved by solver $s$ within $\tau$ seconds and is defined as
@@
We showcase plots of performance profiles and tables of shifted geometric means for different classes of problems supported in the Clarabel solver. We included further detailed numerical results for all solvers in §D, including the iteration counts and the total time.
@@ 4.2 Quadratic programming
where $x\in\textbf{R}^{n}$ (the variable) represents the ratio of allocated assets, $\mu\in\textbf{R}^{n}$ is the vector of expected returns, $\gamma>0$ is the risk-aversion parameter, and $\Sigma\in\mathbf{S}_{+}^{n}$ the risk covariance matrix which is of the form $\Sigma=FF^{\top}+D$ with $F\in\textbf{R}^{p\times n}$ and $D\in\textbf{R}^{p\times p}$ diagonal. We set the rank $p$ to the integer closest to $0.1n$, and vary $n$ from $5000$ to $25000$.
@@
where $a_{i}^{\top}$ is the $i$-th row of $A$ and the Huber loss $\phi_{h}:\textbf{R}\rightarrow\textbf{R}$ is defined as
@@
Figure 2: Performance profiles for the large QPs problem set
(a) Relative performance profile
(b) Absolute performance profile
ClarabelGPU ClarabelRs Gurobi Mosek
Shifted GM Full Acc. 1.0 11.29 2.13 4.18
Low Acc. 1.0 11.29 2.13 4.18
Failure Rate (%) Full Acc. 0.0 20.0 0.0 0.0
Low Acc. 0.0 10.0 0.0 0.0
(c) Benchmark timings as shifted geometric mean and failure rates
Results for large QP tests are shown in Figure 2. We benchmark ten different examples from two classes above and set the time limit to $1$h. We compare our GPU implementation ClarabelGPU with ClarabelRs, the Rust version 0.9.0 version of Clarabel equipped with the faer multithreaded linear system solver; and two commercial solvers, Gurobi and Mosek. ClarabelGPU is the fastest solver on these problems, and it has the lowest per-iteration time for almost all examples. Since most of time of an interior point solver is spent on factorizing and solving a linear system in QPs, we can say that ClarabelGPU benefits from the use of the CUDSS linear system solver and it is more than 2 times faster than Gurobi, about 4 times faster than Mosek and 10x times faster than the existing Rust implementation with the multithreaded faer linear system solver.

### v3

@@ 3.3 Mixed parallel computing strategy
The cone operations related to steps 4, 6, 7, 9, and 10 above require only information local to each constituent cone, and hence can be executed concurrently with respect to individual cones. We propose the mixed parallel computing strategy for a cone operation across different types of cones. Each family of cones is handled in parallel, and families of cones can be addressed by separate streams.
As stated earlier, the cone operations related to zero cones and nonnegative cones are simply vector additions and multiplications, which have already been parallelized in CUDA. In our implementation we aggregate all zero cones into a single zero cone in our preprocessing step. The same holds for nonnegative cones. For the remaining cones, we parallelize within each family of cones (i.e., second-order cone, exponential cone, etc.). For each family of cones, we allocate a thread to each cone belonging to that family, and execute in parallel. This adheres to the SIMD computing paradigm, as each type of cone has its own set of instructions for updating its scaling matrix.
We employ our mixed parallel computing strategy in step 4, the scaling matrix update. Recall that the conic constraint $\mathcal{K}$ can be decomposed as a Cartesian product of $p$ constituent atomic cones $\mathcal{K}_{1}\times\dots\times\mathcal{K}_{p}$ that are ordered by cone types. We assume $\mathcal{K}_{1}$ is a zero cone, $\mathcal{K}_{2}$ is a nonnegative cone, $\mathcal{K}_{3}$ to $\mathcal{K}_{i}$ are second-order cones, $\mathcal{K}_{i+1}$ to $\mathcal{K}_{j}$ are exponential cones, $\mathcal{K}_{j+1}$ to $\mathcal{K}_{l}$ are power cones and $\mathcal{K}_{l+1}$ to $\mathcal{K}_{p}$ are positive semidefinite cones. Due to our decomposition into constituent atomic cones, $H$ is a block-diagonal matrix,
@@
Algorithm 2 Algorithmic sketch for the scaling matrix $H^{k}$ update in Algorithm 1
0: Current iterate $v^{k}$, streams $st_{soc},st_{exp},st_{pow},st_{sdp}$.
1:function _update_H($v^{k}$)
2:   Update $H_{\text{zero}}^{k}$   // Zero cone ($t=1$)
3:   Update $H_{\text{nn}}^{k}$   // Nonnegative cone ($t=2$)
4:
5:   // Second-order cones
[…]
7:
8:   // Exponential cones
[…]
10:
11:   // Power cones
[…]
13:
14:   // Positive semidefinite cones
[…]
16:
17:   // Synchronize scaling update
18:   CuDeviceSynchronize()
19:
20:   return $H^{k}$   // Output the scaling matrix $H^{k}$
21:end
Algorithm 2 illustrates the update of scaling matrix $H^{k}$ in function _update_H($v^{k}$). For the update of $H$ at iteration $k$, we first set diagonal terms of the scaling matrix $H_{1}^{k}$ for the zero cone to all 0s, and then update the diagonals of the scaling matrix $H_{2}^{k}$ for the nonnegative cone by element-wise vector division, as in the Nesterov-Todd (NT) scaling [56]. These computations are already parallelized by CUDA. For conic constraints other than the zero cone and the nonnegative cone, the corresponding parts in the scaling matrix $H$ are no longer diagonal, but we observe $H_{t}$ only requires local information within each constituent cone, and thus we can solve for each $H_{t}$ concurrently, independent to the other cones. We implement kernel functions _kernel_soc_update_H(), _kernel_exp_update_H(), _kernel_pow_update_H() and _kernel_sdp_update_H() for second-order cones, exponential cones, power cones and positive semidefinite cones respectively. Details regarding the update functions for each type of cone can be found in §A. Each kernel function will process one cone per thread, and updating scaling matrices of the same class of cones is executed in parallel. Kernel functions for different cone classes are launched independently in separate streams. Synchronization is required at the end of the scaling matrix update.
@@
Note that the SIMD GPU computing structure naturally favors balanced workloads, i.e., the workloads in each thread should be similar so that the synchronization will not take too much time. The exponential and power cones are 3-dimensional nonsymmetric cones that all cone operations have balanced workload among the same class of cones, thus we can parallelize the computation for each cone. On the other hand, the second-order cones may vary in dimensionality, which may not mesh well with the SIMD computing paradigm. In most use cases for large-scale second-order cone problems, e.g., optimal power flow problems [2] and finite-element problems [5], the second-order cones are of small (less than 5) dimensionality; thus the effect of this workload imbalance is negligible. Regardless, we support second-order cones of all sizes. In the next section, we outline how to process second-order cones of high dimensionality.
3.4 Dynamic parallelism for second-order cones
Due to the definition of the barrier function (6) over a second-order cone $(t,x)\in\mathcal{K}_{q}^{n}$, the computation of residuals involving $t^{2}-\|x\|^{2}$ requires substantial reduction operations within each cone. This introduces an additional level of parallelism, particularly relevant when the dimension of a second-order cone reaches thousands or more, as seen in applications like multistage portfolio optimization.
In such high-dimensional cases, processing second-order cones in a purely thread-wise fashion above becomes inefficient, as the residual computation becomes effectively single-threaded. A straightforward workaround is to use the dot product operations provided by cuBLAS for computing residuals. However, this approach processes each high-dimensional cone sequentially and fails to exploit cone-level parallelism when multiple second-order cones are present. To address this, we implement custom reduction operations and leverage dynamic parallelism in CUDA, which enables additional cone-level parallelism and better utilization of GPU resources.
Algorithm 3 illustrates how to compute $q$ residuals via dynamic parallelism. Suppose we have a vector $x:=[x_{1};\dots;x_{q}]$ where $x_{i}:=(t_{i},u_{i})\in\mathcal{K}_{q}^{n_{i}}$ and residual vector $r\in\textbf{R}^{q}$ that stores the residual of cone $i$, i.e., $r[i]=t_{i}^{2}-\|u_{i}\|^{2}$. We first launch _parent_kernel_soc_residual($x,r$) for the residual computation. We then launch $q$ child kernel functions, _child_kernel_soc_residual ($x_{i},r$), for each thread $i$ ($1\leq i\leq q$) within the parent kernel function, processing the residual of cone $i$. The $2$-norm of $u_{i}$ can be realized by the standard parallel reduction utilizing shared memory efficiently, see Algorithm 4. Finally, we obtain the residual for each cone $i$ and store it back to $r$ in the child function.
Algorithm 3 Dynamic parallelism for the residual computation of $q$ second-order cones
[…]
1:
2: // Parent function
3:function _parent_kernel_soc_residual($x,r$)
4:   $i\leftarrow$ (blockIdx.x -1)* blockDim.x + threadIdx.x  //Julia is 1-indexed
5:   if ($i\leq q$)
6:     $x_{i}\leftarrow$ extract cone block $i$ from $x$
7:
8:     //compute residual of each cone $i$
9:     _child_kernel_soc_residual($x_{i},r$)
10:   end
11:end
12:
13: // Child function
14:function _child_kernel_soc_residual($x_{i},r$)
15:   $t_{i}\leftarrow x_{i}[1]$ // Julia: 1-based indexing
16:   $u_{i}\leftarrow x_{i}[2:end]$
17:   $\|u\|_{2}^{2}\leftarrow\texttt{\_reduction\_norm2}(u_{i},n_{i})$
18:   $r[i]\leftarrow t_{i}^{2}-\|u_{i}\|_{2}^{2}$
19:end
3.5 Batched support for positive semidefinite cones
The positive semidefinite (PSD) cone is a matrix cone that can vary in dimensionality. Computing scaling matrices and step sizes involves several matrix factorizations, such as Cholesky factorization, singular value decomposition and eigenvalue decomposition [52]. In CuClarabel, we address this by using batched matrix factorizations provided by the cuSOLVER library. Hence, we only support a restricted class of SDPs in which all PSD cones have the same dimensionality that is less than or equal to $32$. This includes examples from finite element analysis problems [57, 58], where each SDP constraint encodes localized, element-wise or point-wise conditions —such as stress admissibility or yield criteria—that can be uniformly applied across the mesh.
@@ 3.7 Solving linear systems
Most of the computation time for our interior point method is spent in factorizing the matrix $K$ and in the three backsolve operations in (15). Our GPU implementation also leverages the power of the newly released sparse linear system solver cuDSS [40] for the $LDL^{T}$ factorization and backsolve operations. We implement a self-contained iterative refinement originated from [59] to increase the numerical stability of the backsolve operation.
@@
    Solve for the residual at step $i$: $r_{i}=b-Kx_{i-1}$, where $x_{i-1}$ is our guess for iteration i1i-1. (Full precision)
@@
In all our matrix solves, given matrix $K$ and right-hand side $b$, we factorize the lower precision copy of $K$, and then solve the linear system in this lower precision. To enhance numerical stability during matrix factorization, a regularization term $\delta I$ with
$\delta:=\delta_{s}+\delta_{d}\max_{i}|D_{ii}|,$
is added to $K$, where $\delta_{s}$ is a static regularization and $\delta_{d}\max_{i}|D_{ii}|$ provides dynamic scaling based on the magnitude of the diagonal entries of $D$. To offset the regularization effect and rounding errors from the lower precision in backsolves, we solve in full precision for the other steps, i.e., we compute the residual $r_{i}$ and save the update $\{x_{i}\}$ in full precision. When we change the factorization data type from Float32 to Float64, it reduces down to the standard iterative refinement as employed in standard conic optimization solvers [15, 29].
We apply iterative refinement until either (1) we reach a pre-specified number of max steps $T_{\max}$ for iterative refinement, or (2) the $\ell_{\infty}$ norm of $b-Kx_{i}$ reaches a certain threshold, specifically $\left\|b-Kx_{i}\right\|_{\infty}\leq t_{\text{abs}}+t_{\text{rel}}\left\|b\right\|_{\infty}$. We set both $t_{\text{abs}}$ (absolute tolerance) and $t_{\text{rel}}$ (relative tolerance) to $10^{-12}$, and the maximum number of steps $T_{\max}=10$. We set $\delta_{s}$ to the square root of machine precision, i.e., $\sqrt{\epsilon}\approx 3.45e^{-4}$, for mixed precision and $\delta_{s}=1e^{-8}$ for full precision, and we set $\delta_{d}$ to the square of machine precision, i.e., $\epsilon^{2}$.
Although the mixed precision for the iterative refinement can improve the numerical stability for solving linear system in lower precision, it is to be expected that the mixed precision may take longer time to converge or fail in cases where the matrix $K$ is extremely ill-conditioned. Caution is required when using mixed precision for numerically hard problems, e.g., conic programs with exponential cones. Hence, we provide the mixed precision as an optional choice and use Float64 by default.
@@ 4 Numerical experiments
We have benchmarked our Julia GPU implementation of the Clarabel solver, denoted as ClarabelGPU, against the state-of-the-art commercial interior-point solvers MOSEK [12] and Gurobi [60]. We also include our Rust CPU implementation of Clarabel solver with the 3rd-party multithreaded supernodal LDL factorization method in the faer-rs package [61], denoted as ClarabelRs. Note that MOSEK and Gurobi already utilize multithreaded linear system solvers. We include benchmark results for several classes of problems including quadratic programming (QP), second-order cone programming (SOCP) and exponential cone programming. All benchmarks are performed using the default settings for each solver, with pre-solve disabled where applicable to ensure equivalent problem-solving conditions. No additional iteration limits are imposed beyond each solver’s internal defaults. We set $\epsilon_{\text{feas}}=1e^{-6}$ for all solvers in termination check. All experiments were carried out on a workstation with Intel(R) Xeon(R) w9-3475X CPU @ 4.8 GHz with 256 GB RAM and NVIDIA GeForce RTX 4090 24GB GPU. All benchmarks tests are scripted in Julia and access solver interfaces via JuMP [82]. We use Rust compiler version 1.76.0 and Julia version 1.10.2. Before recording benchmark tests, we run a few example solves; this offsets the just-in-time (JIT) precompilation overhead in Julia.
@@ 4.1 Benchmarking metrics
We choose the same benchmarking tests as used in the Clarabel solver [15], and compare our results using metrics that are commonly used when comparing computational time across different solvers. For a set of $N$ test problems, we define the shifted geometric mean $g_{s}$ as
@@
where $\mathcal{I}_{\leq\tau}(u)=1$ if $\tau\leq u$ and $\mathcal{I}_{\leq\tau}(u)=0$ otherwise. For $\tau=1$, $f^{r}_{s}(1)$ denotes the ratio of problems where the solver $s$ performs the best, and $\sum_{s}f^{r}_{s}(1)$ should be equal to 1 if every problem is solvable for at least one solver. We also compute the absolute performance profile $f^{a}_{s}:\textbf{R}_{+}\mapsto[0,1]$, which denotes the fraction of problems solved by solver $s$ within $\tau$ seconds and is defined as
@@
Our numerical experiments highlight the performance gains achieved by the GPU implementation of Clarabel on a diverse set of conic optimization problems. We record total time by default, the sum of the setup time, including data equilibration and solver initialization, and the solve time, which is the running time for the algorithm underlying a solver.
@@ 4.2 Quadratic programming
where $x\in\textbf{R}^{n}$ (the variable) represents the ratio of allocated assets, $\mu\in\textbf{R}^{n}$ is the vector of expected returns, $\gamma>0$ is the risk-aversion parameter, and $\Sigma\in\mathbf{S}_{+}^{n}$ the risk covariance matrix which is of the form $\Sigma=FF^{T}+D$ with $F\in\textbf{R}^{p\times n}$ and $D\in\textbf{R}^{p\times p}$ diagonal. We set the rank $p$ to the integer closest to $0.1n$, and vary $n$ from $5000$ to $25000$.
@@
where $a_{i}^{T}$ is the $i$-th row of $A$ and the Huber loss $\phi_{h}:\textbf{R}\rightarrow\textbf{R}$ is defined as
@@
Table 1: Benchmarking for quadratic programming
iterations total time (s)
Problem ClarabelGPU Mosek* ClarabelRs Gurobi ClarabelGPU Mosek* ClarabelRs Gurobi
portfolio_optimization_5000 15 9 15 16 0.372 0.936 7.36 0.513
portfolio_optimization_10000 18 10 18 17 1.42 4.24 48.8 2.8
portfolio_optimization_15000 18 11 18 17 3.55 12.7 147 7.52
portfolio_optimization_20000 17 12 17 17 6.66 28.5 307 14.7
portfolio_optimization_25000 18 11 18 18 11.8 49.3 549 27.2
huber_fitting_5000 8 11 8 9 4.81 13.5 65.6 48.2
huber_fitting_10000 8 11 8 9 37.1 55.1 612 276
huber_fitting_15000 8 13 8 10 122 289 2.1e+03 747
huber_fitting_20000 8 13 - 10 283 807 - 1.76e+03
huber_fitting_25000 8 15 - 10 547 1.83e+03 - 1.24e+03
Results for large QP tests are shown in Table LABEL:table:qp. We benchmark ten different examples from two classes above and set the time limit to 11h. We compare our GPU implementation ClarabelGPU with ClarabelRs and two commercial solvers, Gurobi and MOSEK. ClarabelGPU is the fastest solver on these problems, and it has the lowest per-iteration time for almost all examples. Since most of time of an interior point solver is spent on factorizing and solving a linear system in QPs, we can say that ClarabelGPU benefits from the use of the cuDSS linear system solver and it is more than 2 times faster than Gurobi, about 4 times faster than MOSEK and 10x times faster than the existing Rust implementation with the multithreaded faer-rs linear system solver.